OpenStudy (anonymous):

Does -4 equal 2/3(-6) ? This is confusing me, help please

OpenStudy (anonymous):

y=\[2\over3x\] Given points are (-6,-4) Does -4 equal 2/3(-6) ?

OpenStudy (anonymous):

Yes, do you have a scientific calculator? Put in 2/3*-6

OpenStudy (lgbasallote):

wait... is that \[\huge y = \frac 2{3x}\] or \[\huge y = \frac 23x\]

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

@Susan, it's 4.

OpenStudy (lgbasallote):

ahh so if you substitute \[\huge \frac{2}{3(-6)} \implies \frac 2{-18}\] do you think this will equal -4?

OpenStudy (anonymous):

i honestly did it on a normal one and got -5.3

OpenStudy (anonymous):

sry got that first one wrong

OpenStudy (lgbasallote):

what first one?

OpenStudy (anonymous):

The first answer I got from calculator from doing what susan said. Nevermind that.

OpenStudy (lgbasallote):

if you use a calculator... remember to do this \[\huge 2 \div (3 \times -6)\] you will get a different answer if you do \[\huge 2 \div 3 \times -6\]

OpenStudy (anonymous):

Can't you cancel though? 2 over 3(-6) ? Cancel out 3 and -6 ? or.. no?

OpenStudy (lgbasallote):

nope

OpenStudy (anonymous):

lgba, enforcer of order of operations. Keep up the good work.

OpenStudy (lgbasallote):

you cancel when you are dividing -6 and 3...but since you're multiplying them, you can't cancel

OpenStudy (lgbasallote):

@SmoothMath just pointing out the pickiness of a calculator/wolframalpha

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

Honestly Speaking I'm still kind of lost, I skipped the problem and i'm trying to get it again.. left the problem up for about 30mins but no one came here so i just completely ditched it.. sorry if im a little slow..

OpenStudy (lgbasallote):

no worries. so where are you lost?

OpenStudy (anonymous):

From 2 over -18. It's all asking me if it's equal to -4 (Apologies the - sign is sort of blurried off the page.

OpenStudy (anonymous):

)* haha..

OpenStudy (lgbasallote):

so you're confused if 2/-18 is equal to -4 or not?

OpenStudy (anonymous):

yeah.

OpenStudy (lgbasallote):

well you're allowed a calculator right? try inputting \(2 \div -18\) what do you get?

OpenStudy (anonymous):

switch the numbers cause the number is wayy off... it's -9

OpenStudy (lgbasallote):

why switch?

OpenStudy (lgbasallote):

you can't switch just because it doesn't show the answer what you want.... what does the calculator say about \(2 \div -18\)?

OpenStudy (anonymous):

because the answer comes out as -0.111...

OpenStudy (anonymous):

or can it not be done?

OpenStudy (lgbasallote):

that means 2/-18 is not equal to -4

OpenStudy (lgbasallote):

so that means point (-6,-4) is not on the line

OpenStudy (anonymous):

okay, can you help me with another similar one? these confuse me alot..

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

if not thanks :P

OpenStudy (anonymous):

kk thanks, sorry the message delayed there. uhm..

OpenStudy (anonymous):

y= -1/4x ; (-8,2)

OpenStudy (lgbasallote):

is that \[\huge y = -\frac 14x\] or \[\huge y = -\frac 1{4x}\]

OpenStudy (anonymous):

the first one, sorry not good at using the equation thing yet lol

OpenStudy (lgbasallote):

so what's your solution?

OpenStudy (anonymous):

I got -32

OpenStudy (anonymous):

Mainly because i noticed you crossed multiplied so i tried that but this is just getting me confused ..i guess i get lost because i get a little confused when i see the points given

OpenStudy (lgbasallote):

i didn't cross multiply...when did i?

OpenStudy (anonymous):

When i saw you got 2/-18

OpenStudy (lgbasallote):

that was because your x was in the denominator. this time, x is in the numerator. so you do \[\huge -\frac 14 (-8) \implies -\frac { 1 \times 8}{4}\] got it?

OpenStudy (lgbasallote):

that should be \[\huge -\frac {1\times -8}4\]

OpenStudy (anonymous):

i think youre getting confused lol, or is it me? x.x.. -1/4x (the variable x) counts for all -1/4

OpenStudy (lgbasallote):

you are confused

OpenStudy (anonymous):

thought as much

OpenStudy (lgbasallote):

if the x is in the middle, it's counted as multiplying x to the numerator

OpenStudy (anonymous):

oh.. okay.

OpenStudy (lgbasallote):

so now.. solve \[\huge \frac{-1 \times -8}4\] what do you get?

OpenStudy (lgbasallote):

by the way..that's another thing \[\huge -\frac 14 \implies \frac{-1}4\]

OpenStudy (anonymous):

Alright. Well, as i was taught, -1x cant be subtracted from 8 because they're not like terms..?

OpenStudy (lgbasallote):

that's not -1x...that's -1 times -8

OpenStudy (anonymous):

then it's 8.

OpenStudy (lgbasallote):

\[\huge \frac{-1x - 8}4\] ^that's -1x

OpenStudy (lgbasallote):

-1 times -8 is 8 right. now divide it by 4

OpenStudy (anonymous):

2

OpenStudy (anonymous):

y=2 then it matches the given points.

OpenStudy (lgbasallote):

right. so do you think (-8,2) is a point on the line?

OpenStudy (anonymous):

Yes I do. I get it a lot more now..

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

thanks a lot for your help. you even taught me some haha

OpenStudy (lgbasallote):

welcome